I need help...the instructions are Find the distance between P and Q.
P (4,9); Q (9,21)
Can someone help me out? Thank You!

Answer by checkley71(8403)

ON THE X AXIS WE HAVE 9-4=5
ON THE Y AXIS WE HAVE 21-9=12
THESE ARE TWO SIDES OF A RIGHT TRIANGLE WITH THE THIRD SIDE BEING THE HYPOTENUSE THUS
H^2=5^2+12^12
H^2=25+144
H^2=169
X=13 IS THE DISTANCE BETWEEN THE TWO POINTS P & Q (4,9)(9,21)